### v2

# Less is More: Data-Efficient Adaptation for Controllable Text-to-Video Generation

Abstract
Fine-tuning large-scale text-to-video diffusion models to add new generative controls, such as those over physical camera parameters (e.g., shutter speed or aperture), typically requires vast, high-fidelity datasets that are difficult to acquire. In this work, we propose a data-efficient fine-tuning strategy that learns these controls from sparse, low-quality synthetic data. We show that not only does fine-tuning on such simple data enable the desired controls, it actually yields superior results to models fine-tuned on photorealistic “real” data. Beyond demonstrating these results, we provide a framework that justifies this phenomenon both intuitively and quantitatively.
Figure 1: Our “Less is More” framework for data-efficient controllable generation. A T2V backbone, fine-tuned solely on a sparse, low-fidelity synthetic dataset (left), learns to generalize to complex physical controls. This enables precise, high-fidelity manipulation of shutter speed (motion blur), aperture (bokeh), and color temperature during real-world inference (right), driven by a continuous control.
1 Introduction
Recent advances in generative AI using diffusion models have enabled unprecedented levels of quality in video generation The primary foundational models for video generation are text-to-video (T2V) such as [wan2025wan, polyak2024movie], where users describe their desired creations in natural language. Due to limitations of controllability via text, significant effort has been put into creating methods that accept other input modalities, such as images, keyframes, depth maps, bounding boxes, pose skeletons, driver videos, camera trajectories, and so on [wan2025wan, jiang2025vace, cheng2025wan, bai2025recammaster]. However, achieving consistent, reliable, and intuitive fine-grained control over all aspects of the output video still remains a challenge. In this work, we tackle the problem of adding a control mechanism over specific low-dimensional physical or optical properties, such as camera intrinsics, via a simple synthetic data generation and model fine-tuning framework.
A common pattern that has emerged in methods for producing specialized generative video models is starting with a large “foundation” model, trained on huge amounts of video data, and fine-tuning it on a carefully crafted task-specific smaller dataset [polyak2024movie]. Such datasets help the model focus on a particular character identity, artistic style, or specialized effect. The success of such approaches hints at the fact that the initial pre-training equips the model with many useful priors implicit in its latent representation, which can be explicitly “coaxed” out during post-training. Our approach is the first approach that aims to enable conditioning on camera effects (like shutter speed, focal length, etc) in pre-trained video generative models for consistent generation. While the quality of the post-training data is indeed critical, we argue and demonstrate that, surprisingly, having data that are perfectly photorealistic and representative of the output domain can not only be unnecessary but even detrimental for certain specializations.
This paper proposes an approach that allows the use of synthetic data to learn several conditioning effects. We contribute a novel joint training approach that factorizes adaptation: a standard low-rank adapter (LoRA) [hu2022lora] encodes a minimal domain shift, while a disentangled cross-attention adapter learns the conditioning physical effect. Additionally, our paper’s contributions include a formal analysis of why this data-efficient approach succeeds. We demonstrate that, counterintuitively, the success of small-data fine-tuning hinges on the simplicity of the synthetic data. We show that fine-tuning on photorealistic synthetic data, while seemingly a higher-fidelity choice, induces catastrophic forgetting by corrupting the backbone’s pre-trained priors, leading to a “content collapse.” To quantify this, we introduce a new evaluation framework that measures this generative drift and its impact on semantic fidelity. Using this framework, we show that a model trained on simple data retains its generative diversity and high semantic fidelity, while the model trained on complex data suffers a quantifiable and catastrophic collapse. Our work provides both a data-efficient method for controllable video generation and a formal methodology for diagnosing and preventing backbone corruption during adaptation.
2 Related Work
Text-to-Video Generation. Diffusion models [ho2020denoising, sohl2015deep] have become the leading approach for video generation, largely by adapting pre-trained text-to-image models with temporal modules. Early methods like AnimateDiff [guo2023animatediff] showed that adding temporal attention to a frozen T2I U-Net [ronneberger2015u] enables high-quality motion. The field has since moved toward Diffusion Transformers (DiTs) [peebles2023scalable], which underpin state-of-the-art open-source systems such as Hunyuan Video [kong2024hunyuanvideo] and Wan [wan2025wan]. Closed-source models including Sora [OpenAI_Sora_2024], Gen-3 [Runway_Gen3_2024], and Kling [Kuaishou_Kling_2024] further push the frontier with long-form, high-fidelity, physically consistent video. Although our experiments use Wan T2V model, our fine-tuning paradigm and evaluation framework are model-agnostic and apply to any video diffusion backbone.
Controllable Generative Models. Recent work introduces explicit control in generative models through spatial conditions such as depth, masks, sketches, boxes, and segmentation maps [bhat2024loosecontrol, wang2023imagen, voynov2023sketch, li2023gligen, couairon2022diffedit]. ControlNet [zhang2023adding] and T2I-Adapters [mou2024t2i] support such inputs efficiently, with extensions to video focusing on camera pose and trajectory [he2024cameractrl, yu2024viewcrafter, wang2024motionctrl]. Methods like Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto] model physical imaging effects, but their image-based designs fail to generalize to video due to temporal inconsistency. Token-based conditioning via CSaT [fang2024camera] offers implicit control but remains data-heavy. Large-scale, richly annotated video datasets for physical camera parameters are challenging to obtain. Image LoRAs [hu2022lora] provide lightweight adaptation for effects including motion blur, bokeh, restoration, and super-resolution [zhang2025lion, fortes2025bokeh, park2024contribution, sun2025pixel]. Our approach builds on these ideas by injecting scalar physical parameters through a decoupled cross-attention adapter for controllable video synthesis, enabled by targeted synthetic data.
Learning from Synthetic and Sparse Data. The difficulty of large-scale video collection has led to work showing that strong representations can emerge from simple synthetic data [tian2024learning, mayer2018makes]. Procedurally generated videos can rival natural-video pretraining [yu2024learning], and synthetic datasets often serve as effective substitutes for real videos [guo2022learning, kim2022transferable]. Synthetic 3D environments also provide precise labels for tasks like optical flow, tracking, and video depth [dosovitskiy2015flownet, zheng2023pointodyssey, yang2024depth], and recent methods use such data for low-shot motion learning [ren2024customize, wu2024lamp]. These results suggest that motion priors need not rely on complex natural data, motivating our use of sparse, task-specific synthetic supervision to condition T2V models on camera parameters such as shutter speed, aperture, and temperature.
Evaluation and Metrics for Generative Video Existing one-shot fine-tuning methods [wu2023tune] and even zero-shot approaches [geyer2023tokenflow, qi2023fatezero, yang2023rerender] still struggle to generate high-quality, varied videos. Even just quantitatively evaluating and comparing against such such models remains a complex, multi-dimensional problem. Common metrics include the Fréchet Video Distance (FVD) [unterthiner2018towards, unterthiner2019fvd], which measures distributional similarity in a learned feature space for action recognition, and CLIP Score [radford2021learning], which quantifies text–video semantic alignment. More recent works have proposed adaptations for video, such as X-CLIP Score [ni2022expanding], or more robust alternatives like VQA Score [lin2024evaluating], which is less prone to error on complex compositional prompts. shuttleworth2025loravsfinetuningillusion further investigate how LoRA and full fine-tuning alter pre-trained diffusion models by analyzing changes in their weight matrices through spectral decomposition. While benchmarks such as VBench [huang2024vbench]—which we adopt in our evaluation—offer a suite of metrics for assessing temporal consistency, object fidelity, and aesthetic quality, we identify a critical methodological gap: standard metrics fail to quantify the intrinsic complexity of the fine-tuning data. We address this with a novel metric as part of our contribution.
3 Method
Our primary objective is to introduce direct, low-dimensional, continuous control over a specific attribute of the output produced by a T2V generative model. We achieve this by fine-tuning a pretrained T2V model on a carefully assembled dataset, which exhibits sufficient variation along the desired control axis. In order to make this fine-tuning possible, we introduce two architectural modifications to the base model. Our overall training architecture is illustrated in Figure 2, and the supplementary material contains complete pseudocode for our method.
Disentangled Conditioning Module Following recent work in controllable generation [ye2023ip], we employ a dedicated cross-attention module in order to inject the conditioning signal. We normalize the condition (which in our experiments is a scalar value) to obtain $c\in[-1,1]$. This condition is projected to a higher-dimension embedding vector via a small multi-layer perceptron (MLP): $e_{\text{cond}}=\operatorname{MLP}_{\text{cond}}(c)$. The embedded conditional signal $e_{cond}$ is then injected into the model using a dedicated cross-attention mechanism that operates in parallel with the main text cross-attention. We insert this conditional cross-attention adapter into the deepest third of the original model’s transformer blocks. For a given query $q$ from the video latent, the output of the cross-attention block is a linear combination of the text-conditioning $y_{\text{text}}$ and the attribute-conditioning $y_{\text{cond}}$. Backbone LoRA Rather than attempting to construct a dataset for fine-tuning that uniformly and consistently exhibits variation along the desired control axis and matches the natural video distribution of the original T2V model, we make the deliberate choice to construct datasets that satisfy only the former requirement. This decision frees us from having to do careful curation of natural videos or difficult photorealistic rendering but rather allows us to focus on creating simplistic synthetic examples that illustrate the conditioning effect. However, fine-tuning on such out-of-domain datasets inevitably induces content drift. To address this, we employ a joint training strategy where a backbone LoRA is optimized alongside the conditioning module. This approach enforces a “separation of concerns”: the backbone LoRA absorbs the synthetic domain shift, allowing the conditional adapter to focus exclusively on disentangling the physical control signal. Critically, at inference time, we revert this drift by selectively discarding the backbone LoRA weights from all blocks that are not equipped with the conditional adapter. This restores the model’s original generative priors in the majority of the network while preserving the learned control mechanism.
Figure 2: Overview of our controllable generation pipeline. To achieve decoupled control, we encode the scalar condition separately from the text guidance via a parallel cross-attention module. During training (top), we optimize the conditional adapter while actively updating the backbone by injecting LoRA layers into all DiT blocks. During inference (bottom), we discard the LoRA weights from the shallow two-thirds of the transformer blocks, retaining only the conditional adapter and backbone LoRA in the deepest third of the blocks. This selective retention enables high-fidelity physical control while minimizing semantic corruption of the backbone.
4 Dataset Construction
We evaluate our method on three scalar camera parameters—shutter speed, aperture, and color temperature—using synthetic datasets that isolate their respective effects: motion blur, depth of field, and color tone. We adopt a simple, non-photorealistic domain to enable precise procedural control and clean variations along each conditioning axis. To ensure diversity within this controlled setting, we apply two forms of scene randomization.
Control Scalar Randomization. The primary driver of generalization in our method is the low-dimensional adapter (Sec. 3). To complement this and ensure a continuous response, we employ a multi-level stratified sampling strategy rather than relying on fixed discrete values. We partition the normalized range [1,1][-1,1] into $N$ equal-width bins and sample uniformly from within each bin, applying this strategy across several hierarchical layers. This “pyramid” approach ensures that our sparse dataset provides a rich, continuous signal by concentrating sampling density near the center of the control space, preventing overfitting to specific values.
Scene Content Randomization. We design a dataset of geometric primitives that balances semantic simplicity with functional expressiveness, avoiding the confounding complexity of photorealistic data. Each training scene is procedurally generated with a randomized composition of moving shapes. This configuration ensures that the necessary conditions for physical control are visually observable—such as varying trajectories for motion blur or overlapping depth planes for aperture—while eliminating unnecessary semantic details that could induce backbone corruption.
Photorealistic Ablation Dataset. To test our “Less is More” hypothesis, we conduct a controlled ablation comparing data content. We construct a high-fidelity “photorealistic” dataset and a “simple synthetic” counterpart. Each dataset consists of only one scene per physical effect, thereby isolating data complexity as the sole experimental variable. The photorealistic scenes, characterized by high semantic complexity, are as follows:
  • Shutter Speed: We simulate motion blur by applying temporal averaging (frame accumulation) to clips from a high-speed (240-fps) video of a dynamic scene.
  • Aperture: We generate depth-of-field variations by applying BokehMe  [Peng2022BokehMe] to a single image featuring a complex foreground subject and detailed background.
  • Color Temperature: We generate color variations by applying a physics-based Kelvin transformation to a single natural outdoor scene with a broad color palette.
This dataset is used in our experiments (Sec. 6) to demonstrate the risks of backbone corruption.
5 Evaluation Methodology
We propose a two-stage evaluation framework to measure the drift and corruption of the model quantitatively. The first stage, our novel Fast Evaluation Protocol (FEP), introduces lightweight metrics to quantify the magnitude and velocity of backbone changes during fine-tuning. The second stage, the Slow Validation Protocol (SVP), employs established, full-scale metrics to assess the final generative quality and temporal coherence, allowing us to interpret the nature of the changes measured by FEP.
5.1 Stage 1: Fast Evaluation Protocol (FEP)
The FEP is a computationally efficient protocol designed to quantify the magnitude of semantic and distributional shifts in a fine-tuning model relative to its original state. The protocol consists of a single-step denoising pass to generate a minimal 4-frame output from a fixed latent seed. This pass is repeated across a broad, semantically diverse prompt set, with specific details provided in our experimental setup (Sec. 6).
FEP Metrics. We define two metrics that compare the single-step output of an adapted checkpoint against the original (pre-trained) backbone in the CLIP embedding space.
  • Single-Step Fidelity (SSF) Score: The average per-prompt cosine similarity between the original backbone’s and the adapted checkpoint’s embeddings. This measures adherence to the original model’s semantic priors. A score near 1.0 indicates minimal semantic change, while a low score signifies a major shift.
  • Single-Step Fréchet Distance (SS-FD): The Fréchet Distance between the full embedding distributions of the original backbone and the adapted checkpoint. This measures holistic distributional divergence. A high score indicates a significant shift in the model’s output distribution away from the original.
  • Baseline Calculation: To establish a baseline for inherent variance, we compute these metrics by comparing the original T2V backbone against itself, using only different random seeds. This baseline captures minimal, non-deterministic drift. An adapted checkpoint is thus considered to have minimal backbone change if its SSF score and SS-FD score is comparable to this baseline value.
Quantifying Dataset Complexity via Drift Rate. The FEP metrics enable us to quantify dataset complexity by its impact on the model. We define the Distributional Drift Rate ($\mathcal{V}_{\text{drift}}$) as the rate of change in our FEP metrics: $\mathcal{V}_{\text{drift}}=\delta(\text{SS-FD})/\delta(\text{steps})$. A low-complexity dataset (e.g., our 2D primitives) exhibits a low $\mathcal{V}_{\text{drift}}$. Conversely, a high-complexity dataset (e.g., photorealistic scenes) induces a high $\mathcal{V}_{\text{drift}}$. This metric thus provides a quantitative measure of dataset complexity by reflecting the magnitude of change it induces in the backbone.
5.2 Stage 2: Slow Validation Protocol (SVP)
The SVP stage is designed to assess the final generative quality of a model using established, full-scale metrics. It involves a full denoising process to assess both semantic fidelity and perceptual quality.
SVP Metrics. We assess two primary categories of metrics:
  • Semantic Fidelity (Backbone Health): We compute video-text alignment using the X-CLIP Score [ni2022expanding] and the VQA score [lin2024evaluating]. Cross-referencing these scores with the $\mathcal{V}_{\text{drift}}$ from FEP allows us to diagnose the adaptation: a high-velocity drift is benign if final scores are high (successful adaptation), but malignant if they collapse (catastrophic forgetting).
  • Video Quality & Temporal Coherence: We use six metrics from VBench [zhang2024evaluationagent] to evaluate the generated video quality independent of text guidance: subject/background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
Table 1: Quantitative SVP Results. Decoupled (“clean”) and Joint (“dirty”) inference methods against the baseline using our SVP and VBench metrics. The evaluation covers all three controls: Shutter (ST), Aperture (AP), and Temperature (TP).
Metric Baseline ST(dirty) ST(clean) AP (dirty) AP (clean) TP (dirty) TP (clean)
Semantic Fidelity
   X-CLIP Score 25.390 25.295 25.587 25.181 25.595 25.487 25.595
   VQA Score (Qwen-VL) 0.522 0.453 0.521 0.427 0.513 0.550 0.532
Video Quality
   Subject Consistency 0.951 0.939 0.946 0.968 0.951 0.960 0.950
   Background Consistency 0.973 0.963 0.969 0.979 0.973 0.975 0.972
   Motion smoothness 0.988 0.983 0.987 0.994 0.987 0.990 0.989
   Dynamic Degree 0.427 0.688 0.469 0.219 0.438 0.406 0.417
   Aesthetic Quality 0.606 0.576 0.604 0.629 0.610 0.620 0.605
   Imaging Quality 0.618 0.531 0.596 0.596 0.633 0.664 0.623
Figure 3: Qualitative results for Joint Inference. Retaining the full backbone LoRA weights is capable of yielding high-quality generation, supporting the robustness of our data-efficient training.
6 Experiments
To validate our “Less is More” hypothesis and analyze how different data strategies impact backbone fidelity, we conduct two primary groups of experiments. We use the WAN 2.1 [wan2025wan] as our T2V backbone for all fine-tuning and evaluation.
6.1 Experimental Design
  • Group 1: Data Complexity (Syn vs. Real). To isolate the effect of data content, we conduct a controlled one-shot comparison. For each camera parameter, we train two models on minimal data: (1) a model trained on a single synthetic scene and (2) a model trained on a single photorealistic scene (per Sec. 4). From each scene, we sample 7 scalar conditions to create the training data. This ablation is evaluated using our FEP on the 800-prompt VBench benchmark [zhang2024evaluationagent].
  • Group 2: Inference Strategy (Decoupled vs. Joint). To evaluate our proposed inference strategy and also ablation study on, we use our full model trained on the complete pyramid synthetic dataset (Sec. 4) for each camera parameter. Then compare two inference modes: Decoupled (“clean” inference, pruning shallow DiT LoRAs) and Joint (“dirty” inference, retaining all DiT LoRAs). This comparison is evaluated using our SVP on a curated 96-prompt high-motion suite, which involves a full denoising process of 50 steps to generate 49 frames.
6.2 Results and Analysis
Our quantitative results are presented in Fig. 4. and Tab. 1.
6.2.1 Group 1: Synthetic Data Prevents Forgetting
Drift Monitoring (FEP): The results (Fig. 4, Top Row Left) strongly support our “Less is More” hypothesis. The FEP monitoring shows that the model trained on photorealistic “Real” data (dashed line) exhibits a much faster rate of change in its FEP scores per training step compared to the model trained on “Syn” data (solid line). This observation validates that complex, photorealistic data induces a significantly faster and more severe contextual drift in the T2V backbone than our simple, synthetic data.
Corruption Analysis (SVP): The SVP validation (Fig. 4, Top Row Right) confirms the nature of this drift. The “Real” model’s semantic scores (X-CLIP, VQA) collapse, indicating malignant, catastrophic forgetting. Conversely, our “Syn” model’s scores remain on par with the baseline, proving simple synthetic data is superior for avoiding backbone corruption.
6.2.2 Group 2: Decoupled Inference Retains Fidelity
FEP & SVP Analysis: The results (Fig. 4, Bottom Row Right) validate the superiority of our proposed Decoupled (“clean”) inference. The FEP analysis shows it barely impacts the backbone, with drift remaining around the baseline. The SVP analysis confirms this, showing minimal semantic corruption (SVP scores change <<2%). Meanwhile, the Joint (“dirty”) inference, while inferior, shows only a slight, acceptable score reduction. This confirms our training is robust and does not cause catastrophic degradation, remaining capable of high-fidelity controllable generation (Fig. 3).
Video Quality (VBench): As shown in Tab. 1, a key finding is the stability of the VBench video quality metrics. Both our “clean” and “dirty” inference on model training on synthetic dataset maintain scores for Subject Consistency, Background Consistency, Motion Smoothness, and Aesthetic Quality that are nearly identical to the original baseline. This is a crucial result: it demonstrates that our dataset selection and inference design successfully disentangles the new physical control without introducing unwanted artifacts or degrading the backbone’s core generative quality, which is a common failure case.
Figure 4: Ablation studies on data complexity and inference strategy. Top Row: Synthetic vs. Real Data. A one-shot comparison of fine-tuning on our low-fidelity synthetic (“Syn”) versus complex photorealistic (“Real”) data. (Left) FEP monitoring tracks SSF and SS-FD over training steps. (Right) SVP validation bar charts show final X-CLIP and VQA scores. Bottom Row: Decoupled (Clean) vs. Joint (Dirty) Inference. An ablation on our full pyramid-trained model. (Left) FEP monitoring tracks SSF and SS-FD against training steps. (Right) SVP validation bar charts compare final X-CLIP and VQA scores for our proposed ”clean” inference and the ”dirty” Inference.
6.3 Visual Condition Comparison
Apart from the comparisons via training data and inference methods, we also conducted some qualitative comparisons to prior works such as WAN 2.1 [wan2025wan], Bokeh Diffusion [fortes2025bokeh], and [Yuan_2024_GenPhoto]. Fig. 5 compares the condition ability of different workflow, text-based prompting in the T2V backbone (WAN 2.1) exhibits significant semantic entanglement. This is particularly evident in the Shutter Speed test (Row 1), where the baseline either struggles to associate shutter speed changes with motion blur, or generates a lower-fidelity scene by creating a “metalic” door instead of the requested glass one. A similar entanglement occurs in the Temperature test (Row 3), where the baseline misinterprets color temperature as weather, incorrectly altering the scene’s environment (e.g., adding snow) rather than its color tone. Our method overcomes these shortcomings by correctly applying the physical effect of shutter speed (motion blur) without this semantic or fidelity confusion. Furthermore, for Aperture (Row 2), our model generates high-fidelity bokeh effects comparable to specialized, data-intensive baselines such as Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto]. These results validate our central hypothesis: our data-efficient synthetic training achieves robust physical disentanglement via a low-rank adapter, eluding the limitations of entangled text prompts and rivaling the quality of data-heavy methods. The supplementary material (see Appendix F) provides a number of additional examples (as videos) of our method that demonstrate the consistency of on a range of control values.
Figure 5: Qualitative comparison of physical controls. We compare our method against text-based prompting in a T2V backbone (WAN 2.1) and specialized, data-intensive image baselines [Yuan_2024_GenPhoto] and Bokeh Diffusion [fortes2025bokeh].
7 Analysis
Beyond our quantitative metrics (Sec. 5), we develop a suite of data-free analysis techniques. These methods directly analyze the geometric properties of the model’s weights and their functional outputs, removing any dependency on specific data samples at test time. To validate our hypotheses on adaptation and disentanglement, we briefly introduce these techniques here and leave the full implementation details in the supplementary material.
7.1 Hypothesis 1: Low-Fidelity Data Prevents Catastrophic Forgetting
Our first hypothesis is that fine-tuning on our low-fidelity “synthetic” dataset results in a “cleaner” backbone LoRA update. We posit that this preserves the pre-trained model’s general knowledge and avoids the catastrophic forgetting that occurs when training on high-fidelity “real” data.
Method: Backbone Intruder Dimension Check. We apply the “intruder dimension” analysis from  shuttleworth2024lora. We perform an SVD on the original backbone weight $W_{pre}$ and its final adapted version $W_{lora}=W_{pre}+\Delta W_{lora}$.
An “intruder” is defined as a new, high-ranking singular vector in $W_{lora}$ that is dissimilar (cosine similarity $\epsilon$) to all singular vectors in $W_{pre}$. A high intruder count is a mathematical signature of catastrophic forgetting; therefore, a low count signifies a successful, non-destructive adaptation to the backbone’s existing spectral structure.
The full results of this analysis are shown in the supplementary material (Section E.1). The findings, which compare models trained on different datasets, align perfectly with our conclusions in Sec. 6. This alignment validates both our “Less is More” hypothesis and the effectiveness of our spectral analysis tools.
7.2 Hypothesis 2: Low-Rank Conditional Representation
Another hypothesis is that the adapter has learned an efficient, compact representation for the physical effect, which is a sign of good generalization.
Method: Effective Rank Test. The output of the conditional branch, $y_{\text{cond}}$, is a linear combination of the value vectors in $V_{\text{cond}}(c)$. We can therefore analyze the dimensionality of this value tensor directly. We generate $V_{\text{cond}}(c)$ for a strong condition ($c=1$), reshape it into a 2D matrix, and compute its SVD to analyze its singular value spectrum, just as we states in Sec. 3 a successfully factorized model should learn an efficient and compact representation for the conditioning. A physical effect, such as motion blur, should not require a high-dimensional feature space. A low-rank representation is a strong indicator of good generalization, suggesting that the model has learned the “essence” of the effect rather than overfitting to training data artifacts.
(a) Joint Training
(b) Adapter-Only Training
Figure 6: Singular Value Spectrum of the Conditional Signal (y_cond_principal) in Block 27. (a) In our jointly trained model, the conditional signal exhibits a sharp spectral decay with an effective rank of 1, proving the model learned an efficient, low-dimensional representation for the pure physical effect. (b) In contrast, for the adapter-only trained model, the signal is high-rank, with a slow spectral decay that mirrors the complexity of the content signal (y_text). This is definitive evidence that the adapter was forced to memorize the content of the training data.
7.3 Comparative Spectral Analysis of Factorization
To validate our factorization hypothesis, we conduct a comparative spectral analysis on our “Good” (Jointly Trained) checkpoint and a “Bad” (Adapter-Only) checkpoint.
Analysis of “Good” Checkpoint: Our “Good” model demonstrates successful factorization. We analyze the singular value spectrum of its conditional signal ($y_{\text{cond}}$) (Figure 6 a), which reveals a classic “elbow” curve with an effective rank of 1. This result confirms the adapter learned an efficient, low-dimensional representation of the *effect*, as the base LoRA handled the domain’s content. This low-rank structure is the key to generalization.
Analysis of “Bad” Checkpoint: In stark contrast, the adapter-only model fails. Its conditional signal ($y_{cond}$) has a high-rank spectrum that decays slowly, mirroring the content signal ($y_{\text{text}}$) (Figure 6 b). This is the root cause of its failure: trained in isolation, the adapter did not learn the simple effect but instead memorized the complex, high-dimensional content of the synthetic training data.
The “Bulldozer Effect”: We term this failure mode the “Bulldozer Effect”. The “Bad” adapter’s signal is not only high-rank (content memorization) but also has a massive magnitude and is orthogonal to the backbone’s content space. Instead of collaborating, the adapter represents its memorized content in a vacant subspace. During inference, this powerful signal “bulldozes” and replaces the text signal ($y_{\text{text}}$), resulting in output dominated by the training scene features.
8 Conclusion and Future Work
As scaling laws continue to yield bigger and better pre-training of generative video models, post-training strategies will become all the more crucial in order to achieve reliable, intuitive, and controllable models for specific use cases. Our paper offers a new paradigm for thinking about the construction of useful datasets for fine-tuning video models. We show that rather than attempting to gather a dataset that is as realistic as possible, a better strategy may be to construct one that is as disentangled as possible, foregoing realism entirely. We posit that the latent representations learned by foundation models are so expressive and the priors for realism so strong, that the role of fine-tuning is coaxing out existing learned properties, interpolated somewhere in the existing latent space, rather than extrapolating and expanding the domain.
While we show the efficacy of our method on three specific control types, the general framework inspires many directions for future work. Future work could explore a unified model that learns to disentangle multiple parameters (e.g., aperture and temperature) from a single joint control vector. Furthermore, this data-efficient paradigm could be extended to other spatial controls, such as learning depth or pose from simple geometric data, potentially reducing the need for complex, photorealistic datasets in those domains as well.
Appendix A Dataset Details
Pyramid Sampling Strategy. We first describe our efficient sampling strategy, which balances coverage of the control space with the cost of generating synthetic scenes. Enumerating all combinations of control scalars and randomized scene settings would lead to a combinatorial explosion, so we adopt a pyramid generation strategy that varies the number of sampled scalars per layer while keeping the sampling distribution uniform.
Specifically, we divide the control range [1,1][-1,1] into five layers and sample $9+7+5+3+1=25$ scalar values using jittered sampling, ensuring that every point in the interval has equal probability of being selected. For each layer, we generate six scenes using our randomized scene-generation procedure (described later in this section), and each scene is rendered across all scalar values assigned to its layer.
[…]
Shutter Speed. Given a sampled scalar $c\in[-1,1]$ and a randomized scene, we synthesize videos using a simple 2D physics-based renderer. Each scene contains 1–3 colored geometric primitives (circle, square, triangle, star) moving with constant velocities on a uniform background inside a $512\times 512$ canvas, with elastic boundary reflections. We map $c$ to an effective frame rate $\mathrm{FPS}(c)$ via a centered log mapping, and set the exposure duration to $1/\mathrm{FPS}(c)$. For each output frame, we then integrate over the exposure window by temporally averaging 32 analytically predicted sub-frames, advancing the shapes according to their motion model at each sub-step. Crucially, the underlying shape trajectories and backgrounds are shared across all shutter values for a given scene, so the only changing factor along the control axis is the strength of the motion blur.
Aperture. For aperture control, we build a 3D depth-of-field image dataset in Blender. Each scene contains 2–4 rigid objects (cubes, spheres, cylinders, cones, pyramids) placed on a ground plane with a vertical back wall, both sharing a neutral or softly colored material. Objects are assigned distinct colors and arranged at different depths between $d_{\min}$ and $d_{\max}$, forming a clear foreground–midground–background ordering. We always select one foreground object as the focus target and attach the camera’s DoF focus to this object.
Given a sampled scalar $c\in[-1,1]$, we map it to a physical $f$-number using the same centered log mapping used for shutter speed, i.e., $c\mapsto\text{fstop}(c)\in[\text{fstop}_{\text{lo}},\text{fstop}_{\text{hi}}]$. For each scene, we then render multiple conditions by varying only the aperture $\text{fstop}(c)\,$ while keeping camera pose, object geometry, materials, and lighting fixed. Lighting consists of a gray world background plus an optional point light placed on a circle around the scene (either per-scene or per-render), which introduces mild but controlled highlights without changing depth layering. This produces aligned image sets in which the foreground remains in focus while mid- and background objects exhibit depth-dependent defocus changes in accordance with the aperture control scalar.
Temperature We start from sharp 2D scenes and apply a global white-balance transformation. Each base scene consists of 2–4 colored shapes (circles, squares, triangles, stars) randomly placed on a $512\times 512$ canvas with a uniform background color. We first render a single reference image per scene without blur or noise. Then, for each sampled scalar $c\in[-1,1]$, we map $c$ to a target color temperature $K(c)\in[K_{\mathrm{lo}},K_{\mathrm{hi}}]$ using a perceptually uniform Mired-based mapping, and apply a Kelvin-to-RGB white-balance adjustment to the entire image. The transform is defined relative to a reference neutral temperature (e.g., $K_{\mathrm{ref}}=6500\,\mathrm{K}$) and optionally re-normalized to preserve average luminance. This yields condition-aligned image sets in which geometry, layout, and background remain fixed, while only the global color cast smoothly shifts from warm to cool as a function of the temperature control scalar.
Appendix B Algorithmic Details
We provide high-level pseudocode for our model’s core attention block and our data-free evaluation methodologies.
B.1 Jointly Trained Attention Block
Algorithm 1 outlines the forward pass of a “WanAttentionBlock” that has been adapted with our joint training strategy. It details how the base LoRA update is integrated into the backbone’s weights and how the conditional signal from the “FPSCrossAttentionAdapter” is computed and combined with the text-based content signal.
Input :  $x$: Input video latent features
$c_{\text{text}}$: Text context embedding
$c_{\text{cond}}$: Scalar physical condition (e.g., in [-1, 1])
$W_{q},W_{k},W_{v},W_{o}$: Pre-trained cross-attention weights
$\Delta W_{\text{lora}}$: Learned base LoRA update
$\text{Adapter}_{\text{cond}}$: Trained conditional adapter module
Output :  $x_{\text{out}}$: Updated video latent features
1ex/* 1. Form the final adapted backbone weights */
$W_{q}^{\prime}\leftarrow W_{q}+\Delta W_{\text{lora\_q}}$
$W_{k}^{\prime}\leftarrow W_{k}+\Delta W_{\text{lora\_k}}$
$W_{v}^{\prime}\leftarrow W_{v}+\Delta W_{\text{lora\_v}}$
$W_{o}^{\prime}\leftarrow W_{o}+\Delta W_{\text{lora\_o}}$
1ex/* 2. Compute Query, Text Keys, and Text Values */
$q\leftarrow\text{norm}(x\cdot(W_{q}^{\prime})^{T})$
$K_{\text{text}}\leftarrow\text{norm}(c_{\text{text}}\cdot(W_{k}^{\prime})^{T})$
$V_{\text{text}}\leftarrow c_{\text{text}}\cdot(W_{v}^{\prime})^{T}$
1ex/* 3. Compute the text-based content signal */
$y_{\text{text}}\leftarrow\textnormal{{Attention}}(q,K_{\text{text}},V_{\text{text}})$
1ex/* 4. Compute Conditional Keys and Values via Adapter */
$e_{\text{cond}}\leftarrow\textnormal{{MLP}}_{\text{cond}}(c_{\text{cond}})$
$K_{\text{cond}},V_{\text{cond}}\leftarrow\text{Adapter}_{\text{cond}}(e_{\text{cond}})$
1ex/* 5. Compute the conditional signal */
$y_{\text{cond}}\leftarrow\textnormal{{Attention}}(q,K_{\text{cond}},V_{\text{cond}})$
1ex/* 6. Combine signals with a learned gate $g$ */
$y_{\text{combined}}\leftarrow y_{\text{text}}+g\cdot y_{\text{cond}}$
1ex/* 7. Apply final output projection */
$x_{\text{out}}\leftarrow x+y_{\text{combined}}\cdot(W_{o}^{\prime})^{T}$
Algorithm 1 Forward Pass of a Jointly Trained Attention Block
B.2 Backbone Spectral Drift Analysis
To support the examination of Hypothesis 1 in Section 7, we evaluate how fine-tuning modifies the backbone’s attention projections. Algorithm 2 describes the “Backbone Intruder Dimension Check.” This data-free method quantifies catastrophic forgetting by measuring the spectral similarity between the original pre-trained backbone weights and the final adapted backbone weights after joint training.
Input :  $W_{\text{pre}}$: A pre-trained backbone weight matrix
$W_{\text{lora}}$: The final adapted backbone weight matrix ($W_{\text{pre}}+\Delta W_{\text{lora}}$)
$k$: Number of top singular vectors to check (e.g., 64)
$\epsilon$: Intruder similarity threshold (e.g., 0.5)
Output :  $N_{\text{intruders}}$: Count of detected intruder dimensions
1ex/* 1. Compute SVD for both weight matrices */
$U_{\text{pre}},\_,\_\leftarrow\textnormal{{SVD}}(W_{\text{pre}})$
$U_{\text{lora}},\_,\_\leftarrow\textnormal{{SVD}}(W_{\text{lora}})$
[…]
/* 2. Check top-k vectors of the adapted backbone */
for $j\leftarrow 1$ to $k$ do
 $u_{\text{lora\_j}}\leftarrow j$-th column of $U_{\text{lora}}$
  1ex/* 3. Find max similarity to any pre-trained vector */
 $S_{\text{max}}\leftarrow 0$
 for $i\leftarrow 1$ to number of columns in $U_{\text{pre}}$ do
    $u_{\text{pre\_i}}\leftarrow i$-th column of $U_{\text{pre}}$
    $s\leftarrow\textnormal{{abs}}(\text{cosine\_similarity}(u_{\text{lora\_j}},u_{\text{pre\_i}}))$
    $S_{\text{max}}\leftarrow\textnormal{{max}}(S_{\text{max}},s)$
  1ex/* 4. Count as intruder if similarity is below threshold */
 if $S_{\text{max}}<\epsilon$ then
    $N_{\text{intruders}}\leftarrow N_{\text{intruders}}+1$
1exreturn $N_{intruders}$
Algorithm 2 Backbone Intruder Dimension Check
B.3 Data-Free Spectral Evaluation of the Conditional Adapter
To complement the comparative spectral analysis of Hypothesis 2 in Section 7, we employ a data-free procedure (“Principal Component Showdown”) summarized in Algorithm 3. This method constructs a diagnostic testbench directly from the model’s weights by extracting the dominant principal directions of the backbone’s attention projections. We then evaluate the conditional adapter’s response to these canonical content directions and compute the singular value spectrum of the resulting signal.
Input :  $W^{\prime}_{q},W^{\prime}_{k},W^{\prime}_{v}$: final adapted backbone projection weights
$\text{Adapter}_{\text{cond}}$: trained conditional adapter module
$c_{\text{strong}}$: a strong condition value (e.g., $c=1.0$)
NN: number of principal components to test (e.g., $N=64$)
Output :  $S_{\text{text}},S_{\text{cond}}$: normalized singular value spectra
$\mathcal{R}_{\text{text}},\mathcal{R}_{\text{cond}}$: effective ranks of backbone and conditional signals
1ex
/* 1. Extract principal content primitives from the backbone */
$U_{q},\_,\_\leftarrow\textnormal{{SVD}}((W^{\prime}_{q})^{T})$
$U_{k},\_,\_\leftarrow\textnormal{{SVD}}((W^{\prime}_{k})^{T})$
$U_{v},\_,\_\leftarrow\textnormal{{SVD}}((W^{\prime}_{v})^{T})$
$q_{\text{test}}\leftarrow\textnormal{{TopN}}(U_{q},N)$
$k_{\text{text\_test}}\leftarrow\textnormal{{TopN}}(U_{k},N)$
$v_{\text{text\_test}}\leftarrow\textnormal{{TopN}}(U_{v},N)$
1ex
/* 2. Simulate backbone response to principal content directions */
$y_{\text{text\_principal}}\leftarrow\textnormal{{Attention}}(q_{\text{test}},k_{\text{text\_test}},v_{\text{text\_test}})$
1ex
/* 3. Generate and simulate conditional response */
$K_{\text{cond}},V_{\text{cond}}\leftarrow\text{Adapter}_{\text{cond}}(c_{\text{strong}})$
$y_{\text{cond\_principal}}\leftarrow\textnormal{{Attention}}(q_{\text{test}},K_{\text{cond}},V_{\text{cond}})$
1ex
/* 4. Compute singular value spectra */
$S_{\text{text}},\_\leftarrow\textnormal{{SVD}}(\text{flatten}(y_{\text{text\_principal}}))$
$S_{\text{cond}},\_\leftarrow\textnormal{{SVD}}(\text{flatten}(y_{\text{cond\_principal}}))$
Normalize $S_{\text{text}},S_{\text{cond}}$ by their leading singular values
1ex
/* 5. Estimate effective ranks */
$\mathcal{R}_{\text{text}}\leftarrow\textnormal{{EffectiveRank}}(S_{\text{text}})$
$\mathcal{R}_{\text{cond}}\leftarrow\textnormal{{EffectiveRank}}(S_{\text{cond}})$
1ex
return $S_{\text{text}},S_{\text{cond}},\mathcal{R}_{\text{text}},\mathcal{R}_{\text{cond}}$
Algorithm 3 Data-Free Spectral Rank Analysis (“Principal Component Showdown”)
Appendix C Training Details
All models are trained on NVIDIA A100 80GB (A100E) GPUs using Wan2.1-T2V-14B as the backbone with bfloat16 precision. Unless noted otherwise, we follow the same configuration across all experiments. Each run trains for up to 1000 epochs with a micro-batch size of 1 per GPU, gradient accumulation of 4, and a single pipeline stage. This corresponds to a global batch size of 4 and yields roughly 18,000 training steps for our 30 shots synthetic datasets. Under this setup, each pyramid training stage completes in approximately one day on two A100E.
We use four condition-adapter tokens, an adapter embedding dimension of 256, and LoRA rank 32 for both the condition adapters and the backbone LoRA, with the adapter gate fixed at 0.5 throughout training. A learning rate of $2{\times}10^{-5}$ with a 100-step warmup is applied. Optimization is performed using the AdamW optimizer[loshchilov2017decoupled] with betas $(0.9,0.99)$, weight decay 0.01, and $\varepsilon=10^{-8}$. All LoRA and adapter parameters are trained in bfloat16 with activation checkpointing enabled for memory efficiency.
Appendix D Experiment Details
In this section, we provide additional details for the experimental comparisons introduced in Section 6.
D.1 Group 1: Data Complexity (Syn. vs. Real)
Figure 7: Comparison of the photorealistic and synthetic one-shot datasets used in Group 1. For each physical effect (shutter speed, aperture, temperature), we show the two extreme control values (c=1c=-1 and $c=1$). Although only the endpoints are visualized, each dataset contains seven aligned conditions in total within the normalized range $c\in[-1,1]$.
Dataset Construction. For the Group 1 comparison, we adopt a controlled one-shot setting in which each physical effect (shutter speed, aperture, temperature) is represented by a single scene containing seven scalar conditions that are aligned across the synthetic and photorealistic datasets (see Figure 7). All conditions are normalized to the shared range $c\in[-1,1]$, ensuring that both datasets express identical control values despite differences in rendering styles.
Evaluation Prompt Details. Following the prompt construction strategy used in VBench—which we also rely on for the FEP metric—we generate a diverse set of 96 motion-centric T2V prompts using an LLM (Gemini 2.5). The prompts span eight categories (animals, architecture, food, humans, lifestyle, plants, scenery, and vehicles), with 12 prompts per category. We use this set to evaluate SVP under a broad range of motion patterns and scene types.
Quantitative Results Recap. As shown in Figure 4, models fine-tuned on our simple synthetic dataset exhibit substantially smaller drift from the pretrained backbone when monitored using the FEP metrics (Section 5.1). In contrast, training on the photorealistic dataset leads to noticeably larger and faster deviation throughout training. This difference in drift is reflected in the final SVP results (Section 5.2): the model trained on synthetic data maintains near-baseline semantic fidelity, while the model trained on photorealistic data shows clear backbone degradation. These findings highlight that, for learning physical controls, low-complexity synthetic data provides a more stable and robust supervision signal.
Figure 8: Visual comparison of backbone corruption during training. The synthetic-trained model remains stable and continues to follow the input prompt, whereas the model trained on photorealistic data rapidly drifts toward copying its training scene (see Fig. 7). This manifests as semantic drift, texture collapse, and global color shifts that intensify over training.
Visualization Evaluation. Beyond the quantitative trends, the backbone degradation becomes most apparent when inspecting the model’s inference outputs during training (see Figure 8). When trained on our synthetic dataset, the model continues to follow the input prompt and preserves the overall appearance of the generated scene. In contrast, the model trained on the photorealistic dataset begins to visually drift toward the training images themselves: details, textures, and global color tones increasingly resemble the underlying photorealistic scene rather than the prompted content. This “copying” behavior appears early in training and progressively intensifies, manifesting as semantic drift, texture collapse, and color hallucination. These visual failures provide an intuitive counterpart to the backbone corruption measured by our FEP and SVP metrics.
D.2 Group 2: Inference Strategy (Decoupled vs. Joint)
While Group 1 highlights the stability advantage of synthetic data during training, here we analyze how inference strategy further affects the final output quality. Specifically, we compare the joint inference procedure (“dirty inference”), which directly passes the adapter-modulated activations into the backbone, against our proposed decoupled inference method (“clean inference”) method, which prunes the shallow DiT blocks, removing the residual content drift before decoding.
Quantitative Results Recap. Similar to the quantitative trends observed in Group 1, the metrics in Group 2 also reveal a clear advantage for our clean inference strategy. As summarized in Table 1 and the corresponding FEP/SVP curves and charts in Figure 4, clean inference consistently yields lower drift and higher semantic fidelity compared to dirty inference across all physical controls. Even when the model is trained on the synthetic dataset—where backbone corruption is already minimal—clean inference further reduces the residual deviations and restores the model’s output closer to the original WAN 2.1 behavior. These quantitative results motivate a closer look at the visual differences between the two inference modes.
Visual Evidence. The visual comparisons further reinforce the quantitative trends. As shown in the last column of Figure 8, applying clean inference to a corrupted checkpoint trained on photorealistic data dramatically restores prompt fidelity and global appearance: the model no longer copies textures or colors from the training scene, and semantic drift is substantially reduced. This demonstrates that much of the degradation observed during training is not permanently baked into the weights but instead arises from residual content leakage that clean inference effectively removes.
Figure 9: Effect of clean vs. dirty inference on a model trained with the synthetic pyramid dataset. Although synthetic training induces only mild backbone drift, dirty inference still injects subtle synthetic-style biases (e.g., flattened shading and softened textures) into the outputs. Applying clean inference removes these artifacts and restores the visual style of the original WAN 2.1 model across different control values.
A similar pattern appears even when the model is trained on the synthetic dataset. Although backbone corruption is minimal in this setting, and the model is capable of yielding high-quality output as shown in Figure 3, dirty inference still injects subtle synthetic-style bias—such as flattened textures or blurred shading—into the final output. As shown in Figure 9, clean inference suppresses these artifacts and produces videos that closely match the visual style and sharpness of the original WAN 2.1 backbone across different control values.
Figure 10: Visual comparison of backbone corruption during training. The synthetic-trained model remains stable and continues to follow the input prompt, whereas the model trained on photorealistic data rapidly drifts toward copying its training scene (see Fig. 7). This manifests as semantic drift, texture collapse, and global color shifts that intensify over training.
Limitations of Real Data. Although clean inference substantially reduces the corruption introduced by real-data training, it does not fully eliminate it—especially under extreme control values. As shown in Figure 10, the model trained on the real shutter speed dataset still exhibits noise and structural artifacts when the control approaches the boundaries (e.g., c=1.0c=-1.0), even after applying clean inference. This occurs because the real sequence contains complex high-entropy motion: articulated limbs, nonlinear trajectories, and strong background parallax. In contrast, the synthetic shutter speed dataset is constructed from simple rigid motion with clean parametric trajectories, providing a much more disentangled supervision signal.
These observations reinforce our Less is More thesis: low-complexity synthetic data is inherently easier for the model to factorize and scales reliably across the full control range, whereas real data introduces irreducible entanglement that clean inference can mitigate but not fully correct. Thus, synthetic data remains the preferred choice for learning stable, well-behaved physical controls.
Appendix E Analysis Details
E.1 Hypothesis 1: The Spectrum of Backbone Adaptation
(a) Similarity Heatmap Comparison
(b) Intruder Dimension Comparison
Figure 11: Backbone content drift across depth for the shutter speed condition. Analysis performed on the value projection $W_{v}$. (a) Max cosine similarity between content directions and FPS-conditioned features. (b) Number of content intruders (cosine >0.5>0.5).
Expanding on Hypothesis 1 in Section 7, we analyze how the complexity of fine-tuning data influences the backbone’s spectral drift under conditional adaptation. Unless otherwise noted, all measurements are performed on the value projection $W_{v}$, though we observe qualitatively similar trends in $W_{q},W_{k},$ and $W_{o}$. Figure 11 reports the results for the shutter speed condition using both real and synthetic training data as discussed in Appendix D. Other physical controls exhibit the same overall behavior: simple synthetic data preserves the pretrained spectral structure, whereas complex real data drives substantial high-rank intrusions into the backbone.
Clean Adaptation (Single Synthetic Scene).
As shown in Fig. 11(a), synthetic shutter speed conditioning results in minimal drift throughout the network. The similarity heatmap remains uniformly low (green), indicating that the adapter modifies $W_{v}$ through small non-destructive adjustments. This corresponds to a clean and stable form of adaptation: the model adjusts to the new condition while largely preserving the pretrained spectral structure. However, the specialization remains narrow, as the synthetic data contains only one simple scene.
Catastrophic Forgetting (Single Complex Scene).
In contrast, Fig. 11(b) reveals that real shutter conditioning induces substantial spectral drift in $W_{v}$, particularly in shallow and mid-depth blocks. A large number of high-similarity intruders (cosine similarity >0.5>0.5) appear as early as blocks 0–12 and accumulate into dense clusters in deeper layers. This is the hallmark of catastrophic forgetting: the model learns strong, scene-specific features that overwrite general-purpose representations. These intruder directions propagate into the conditioned path during inference, producing the “ghosting” artifacts seen in the qualitative comparisons. Similar behavior appears under other real-data conditions.
Overall, this data-free analysis corroborates our quantitative findings and highlights that deeper blocks inherently provide cleaner, more stable channels for condition modulation.
Figure 12: Hypothesis 2: Adapter-only training entangles condition and context. On the one-shot temperature-conditioned synthetic dataset, both models learn some temperature effect, but the adapter-only model introduces unintended scene drift, whereas joint training yields clean, isolated temperature changes.
E.2 Hypothesis 2: Adapter-Only Training Entangles Context and Condition
In Hypothesis 2 from Section 7, our comparative spectral analysis showed that the “Good” (jointly trained) checkpoint produces a low-rank, effect-specific conditional signal, whereas the “Bad” (adapter-only) checkpoint produces a high-rank signal that mixes effect and training-scene content. To visualize how this structural difference affects behavior, we evaluate both models on the same one-shot temperature-conditioned synthetic dataset used during training.
Although the dataset is extremely small (1 scene and 7 conditions), both models still learn the temperature effect to a noticeable degree. The purpose of this experiment is not to evaluate the realism of temperature control, but to examine whether each training strategy can express the effect cleanly, without unintentionally altering unrelated scene content.
Adapter-Only Training (High-Rank Conditional Signal).
As shown in Fig. 12, varying the temperature condition $c\in[-1,1]$ causes the adapter-only model to drift away from the original scene: color tone changes are partially entangled with shifts in background structure and global appearance. While the temperature effect itself is present, it is accompanied by unintended modifications to the scene. This matches our spectral findings: the high-rank conditional signal encodes both the intended effect and the training-scene context, which becomes increasingly dominant as $|c|$ increases.
Joint Backbone and Adapter Training (Low-Rank Conditional Signal).
In contrast, the jointly trained model expresses the same learned temperature effect more cleanly. As $c$ varies, the scene content remains stable, and most changes are confined to the desired color-temperature direction. Despite the one-shot nature of the dataset, the jointly trained model avoids leaking training-scene content into the conditional pathway. This aligns with the low-rank spectrum observed in Figure 6, where the effect is captured in a compact direction that interacts with the backbone without re-encoding context.
Overall, even with limited one-shot synthetic data, this diagnostic highlights a consistent pattern: joint training preserves context and isolates the intended effect, whereas adapter-only training entangles the two.
Appendix F Example Gallery
Figure 13 presents a diverse set of examples generated with our shutter-, aperture-, and temperature-controlled T2V framework. Each physical control is parameterized within a unified range $c\in[-1,1]$, enabling smooth and interpretable adjustments to motion blur, depth of field, and color tone. Across a wide variety of scenes and prompts, the outputs change in a consistent and approximately monotonic manner as $c$ varies, demonstrating the reliability and generality of our learned control space.
To further assess this generalization, Figures 14,  15 and 17 evaluate each control dimension in more challenging scenarios. Shutter control extends to moving-camera settings and scenes with multiple independently moving objects, while maintaining predictable blur variation. Aperture control remains stable across diverse depth layouts and can shift focus to locations beyond the foreground, producing smooth bokeh shifts as depth varies (see Figure 16). Temperature control remains consistent across indoor environments and highly stylized domains such as anime and pixel art. Together, these results show that our learned controls generalize far beyond the simple one-shot synthetic scenes used for training.
Figure 13: Qualitative results of our controllable generation. Our model demonstrates precise and continuous control over shutter speed (Row 1-2, motion blur), aperture (Row 3-4, bokeh), and color temperature (Row 5-6) by varying the conditional input $c$ from -1.0 to 1.0 across diverse, high-fidelity video prompts.
Figure 14: Generalization of shutter control to scenes with complex motion. The model responds reliably to the shutter scalar in settings involving moving cameras (e.g., camera-follow and first-person views) and scenes with multiple independently moving objects.
Figure 15: Generalization of aperture control across diverse depth layouts and focal targets. The model handles scenes with multiple depth layers and varied spatial arrangements, and can focus on locations beyond the foreground (e.g., mid- or background planes).
Figure 16: Despite being trained only on images, the model renders smooth bokeh variation as depth changes, enabled by the backbone’s strong prior.
Figure 17: Generalization of temperature control across a wide range of scene types. The model produces stable cooler-to-warmer transitions in indoor environments as well as highly stylized domains such as anime and pixel art.

### v3

@@
 1 Introduction
-Recent advances in generative AI using diffusion models have enabled unprecedented levels of quality in video generation The primary foundational models for video generation are text-to-video (T2V) such as [wan2025wan, polyak2024movie], where users describe their desired creations in natural language.
-Due to limitations of controllability via text, significant effort has been put into creating methods that accept other input modalities, such as images, keyframes, depth maps, bounding boxes, pose skeletons, driver videos, camera trajectories, and so on [wan2025wan, jiang2025vace, cheng2025wan, bai2025recammaster].
+Recent advances in generative AI using diffusion models have enabled unprecedented levels of quality in video generation The primary foundational models for video generation are text-to-video (T2V) such as [43, 28], where users describe their desired creations in natural language.
+Due to limitations of controllability via text, significant effort has been put into creating methods that accept other input modalities, such as images, keyframes, depth maps, bounding boxes, pose skeletons, driver videos, camera trajectories, and so on [43, 15, 3, 1].
 However, achieving consistent, reliable, and intuitive fine-grained control over all aspects of the output video still remains a challenge.
 In this work, we tackle the problem of adding a control mechanism over specific low-dimensional physical or optical properties, such as camera intrinsics, via a simple synthetic data generation and model fine-tuning framework.
-A common pattern that has emerged in methods for producing specialized generative video models is starting with a large “foundation” model, trained on huge amounts of video data, and fine-tuning it on a carefully crafted task-specific smaller dataset [polyak2024movie].
+A common pattern that has emerged in methods for producing specialized generative video models is starting with a large “foundation” model, trained on huge amounts of video data, and fine-tuning it on a carefully crafted task-specific smaller dataset [28].
 Such datasets help the model focus on a particular character identity, artistic style, or specialized effect.
@@
 This paper proposes an approach that allows the use of synthetic data to learn several conditioning effects.
-We contribute a novel joint training approach that factorizes adaptation: a standard low-rank adapter (LoRA) [hu2022lora] encodes a minimal domain shift, while a disentangled cross-attention adapter learns the conditioning physical effect.
+We contribute a novel joint training approach that factorizes adaptation: a standard low-rank adapter (LoRA) [13] encodes a minimal domain shift, while a disentangled cross-attention adapter learns the conditioning physical effect.
 Additionally, our paper’s contributions include a formal analysis of why this data-efficient approach succeeds.
@@
 Text-to-Video Generation.
-Diffusion models [ho2020denoising, sohl2015deep] have become the leading approach for video generation, largely by adapting pre-trained text-to-image models with temporal modules.
-Early methods like AnimateDiff [guo2023animatediff] showed that adding temporal attention to a frozen T2I U-Net [ronneberger2015u] enables high-quality motion.
-The field has since moved toward Diffusion Transformers (DiTs) [peebles2023scalable], which underpin state-of-the-art open-source systems such as Hunyuan Video [kong2024hunyuanvideo] and Wan [wan2025wan].
-Closed-source models including Sora [OpenAI_Sora_2024], Gen-3 [Runway_Gen3_2024], and Kling [Kuaishou_Kling_2024] further push the frontier with long-form, high-fidelity, physically consistent video.
+Diffusion models [12, 36] have become the leading approach for video generation, largely by adapting pre-trained text-to-image models with temporal modules.
+Early methods like AnimateDiff [10] showed that adding temporal attention to a frozen T2I U-Net [32] enables high-quality motion.
+The field has since moved toward Diffusion Transformers (DiTs) [26], which underpin state-of-the-art open-source systems such as Hunyuan Video [17] and Wan [43].
+Closed-source models including Sora [24], Gen-3 [33], and Kling [38] further push the frontier with long-form, high-fidelity, physically consistent video.
 Although our experiments use Wan T2V model, our fine-tuning paradigm and evaluation framework are model-agnostic and apply to any video diffusion backbone.
 Controllable Generative Models.
-Recent work introduces explicit control in generative models through spatial conditions such as depth, masks, sketches, boxes, and segmentation maps [bhat2024loosecontrol, wang2023imagen, voynov2023sketch, li2023gligen, couairon2022diffedit].
-ControlNet [zhang2023adding] and T2I-Adapters [mou2024t2i] support such inputs efficiently, with extensions to video focusing on camera pose and trajectory [he2024cameractrl, yu2024viewcrafter, wang2024motionctrl].
-Methods like Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto] model physical imaging effects, but their image-based designs fail to generalize to video due to temporal inconsistency.
-Token-based conditioning via CSaT [fang2024camera] offers implicit control but remains data-heavy.
+Recent work introduces explicit control in generative models through spatial conditions such as depth, masks, sketches, boxes, and segmentation maps [2, 44, 42, 18, 4].
+ControlNet [55] and T2I-Adapters [22] support such inputs efficiently, with extensions to video focusing on camera pose and trajectory [11, 51, 45].
+Methods like Bokeh Diffusion [7] and Generative Photography [53] model physical imaging effects, but their image-based designs fail to generalize to video due to temporal inconsistency.
+Token-based conditioning via CSaT [6] offers implicit control but remains data-heavy.
 Large-scale, richly annotated video datasets for physical camera parameters are challenging to obtain.
-Image LoRAs [hu2022lora] provide lightweight adaptation for effects including motion blur, bokeh, restoration, and super-resolution [zhang2025lion, fortes2025bokeh, park2024contribution, sun2025pixel].
+Image LoRAs [13] provide lightweight adaptation for effects including motion blur, bokeh, restoration, and super-resolution [56, 7, 25, 37].
 Our approach builds on these ideas by injecting scalar physical parameters through a decoupled cross-attention adapter for controllable video synthesis, enabled by targeted synthetic data.
 Learning from Synthetic and Sparse Data.
-The difficulty of large-scale video collection has led to work showing that strong representations can emerge from simple synthetic data [tian2024learning, mayer2018makes].
-Procedurally generated videos can rival natural-video pretraining [yu2024learning], and synthetic datasets often serve as effective substitutes for real videos [guo2022learning, kim2022transferable].
-Synthetic 3D environments also provide precise labels for tasks like optical flow, tracking, and video depth [dosovitskiy2015flownet, zheng2023pointodyssey, yang2024depth], and recent methods use such data for low-shot motion learning [ren2024customize, wu2024lamp].
+The difficulty of large-scale video collection has led to work showing that strong representations can emerge from simple synthetic data [39, 21].
+Procedurally generated videos can rival natural-video pretraining [52], and synthetic datasets often serve as effective substitutes for real videos [9, 16].
+Synthetic 3D environments also provide precise labels for tasks like optical flow, tracking, and video depth [5, 57, 48], and recent methods use such data for low-shot motion learning [31, 47].
 These results suggest that motion priors need not rely on complex natural data, motivating our use of sparse, task-specific synthetic supervision to condition T2V models on camera parameters such as shutter speed, aperture, and temperature.
-Evaluation and Metrics for Generative Video Existing one-shot fine-tuning methods [wu2023tune] and even zero-shot approaches [geyer2023tokenflow, qi2023fatezero, yang2023rerender] still struggle to generate high-quality, varied videos.
+Evaluation and Metrics for Generative Video Existing one-shot fine-tuning methods [46] and even zero-shot approaches [8, 29, 49] still struggle to generate high-quality, varied videos.
 Even just quantitatively evaluating and comparing against such such models remains a complex, multi-dimensional problem.
-Common metrics include the Fréchet Video Distance (FVD) [unterthiner2018towards, unterthiner2019fvd], which measures distributional similarity in a learned feature space for action recognition, and CLIP Score [radford2021learning], which quantifies text–video semantic alignment.
-More recent works have proposed adaptations for video, such as X-CLIP Score [ni2022expanding], or more robust alternatives like VQA Score [lin2024evaluating], which is less prone to error on complex compositional prompts.
-shuttleworth2025loravsfinetuningillusion further investigate how LoRA and full fine-tuning alter pre-trained diffusion models by analyzing changes in their weight matrices through spectral decomposition.
-While benchmarks such as VBench [huang2024vbench]—which we adopt in our evaluation—offer a suite of metrics for assessing temporal consistency, object fidelity, and aesthetic quality, we identify a critical methodological gap: standard metrics fail to quantify the intrinsic complexity of the fine-tuning data.
+Common metrics include the Fréchet Video Distance (FVD) [40, 41], which measures distributional similarity in a learned feature space for action recognition, and CLIP Score [30], which quantifies text–video semantic alignment.
+More recent works have proposed adaptations for video, such as X-CLIP Score [23], or more robust alternatives like VQA Score [19], which is less prone to error on complex compositional prompts.
+Shuttleworth et al.
+[35] further investigate how LoRA and full fine-tuning alter pre-trained diffusion models by analyzing changes in their weight matrices through spectral decomposition.
+While benchmarks such as VBench [14]—which we adopt in our evaluation—offer a suite of metrics for assessing temporal consistency, object fidelity, and aesthetic quality, we identify a critical methodological gap: standard metrics fail to quantify the intrinsic complexity of the fine-tuning data.
 We address this with a novel metric as part of our contribution.
@@
 Our overall training architecture is illustrated in Figure 2, and the supplementary material contains complete pseudocode for our method.
-Disentangled Conditioning Module Following recent work in controllable generation [ye2023ip], we employ a dedicated cross-attention module in order to inject the conditioning signal.
+Disentangled Conditioning Module Following recent work in controllable generation [50], we employ a dedicated cross-attention module in order to inject the conditioning signal.
 We normalize the condition (which in our experiments is a scalar value) to obtain $c\in[-1,1]$.
@@
   • Shutter Speed: We simulate motion blur by applying temporal averaging (frame accumulation) to clips from a high-speed (240-fps) video of a dynamic scene.
-  • Aperture: We generate depth-of-field variations by applying BokehMe  [Peng2022BokehMe] to a single image featuring a complex foreground subject and detailed background.
+  • Aperture: We generate depth-of-field variations by applying BokehMe  [27] to a single image featuring a complex foreground subject and detailed background.
   • Color Temperature: We generate color variations by applying a physics-based Kelvin transformation to a single natural outdoor scene with a broad color palette.
@@
 SVP Metrics. We assess two primary categories of metrics:
-  • Semantic Fidelity (Backbone Health): We compute video-text alignment using the X-CLIP Score [ni2022expanding] and the VQA score [lin2024evaluating].
+  • Semantic Fidelity (Backbone Health): We compute video-text alignment using the X-CLIP Score [23] and the VQA score [19].
 Cross-referencing these scores with the $\mathcal{V}_{\text{drift}}$ from FEP allows us to diagnose the adaptation: a high-velocity drift is benign if final scores are high (successful adaptation), but malignant if they collapse (catastrophic forgetting).
-  • Video Quality & Temporal Coherence: We use six metrics from VBench [zhang2024evaluationagent] to evaluate the generated video quality independent of text guidance: subject/background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
+  • Video Quality & Temporal Coherence: We use six metrics from VBench [54] to evaluate the generated video quality independent of text guidance: subject/background consistency, motion smoothness, dynamic degree, aesthetic quality, and imaging quality.
 Table 1: Quantitative SVP Results.
@@
 To validate our “Less is More” hypothesis and analyze how different data strategies impact backbone fidelity, we conduct two primary groups of experiments.
-We use the WAN 2.1 [wan2025wan] as our T2V backbone for all fine-tuning and evaluation.
+We use the WAN 2.1 [43] as our T2V backbone for all fine-tuning and evaluation.
 6.1 Experimental Design
@@
 From each scene, we sample 7 scalar conditions to create the training data.
-This ablation is evaluated using our FEP on the 800-prompt VBench benchmark [zhang2024evaluationagent].
+This ablation is evaluated using our FEP on the 800-prompt VBench benchmark [54].
   • Group 2: Inference Strategy (Decoupled vs.
@@
 6.3 Visual Condition Comparison
-Apart from the comparisons via training data and inference methods, we also conducted some qualitative comparisons to prior works such as WAN 2.1 [wan2025wan], Bokeh Diffusion [fortes2025bokeh], and [Yuan_2024_GenPhoto].
+Apart from the comparisons via training data and inference methods, we also conducted some qualitative comparisons to prior works such as WAN 2.1 [43], Bokeh Diffusion [7], and [53].
 Fig.
@@
 Our method overcomes these shortcomings by correctly applying the physical effect of shutter speed (motion blur) without this semantic or fidelity confusion.
-Furthermore, for Aperture (Row 2), our model generates high-fidelity bokeh effects comparable to specialized, data-intensive baselines such as Bokeh Diffusion [fortes2025bokeh] and Generative Photography [Yuan_2024_GenPhoto].
+Furthermore, for Aperture (Row 2), our model generates high-fidelity bokeh effects comparable to specialized, data-intensive baselines such as Bokeh Diffusion [7] and Generative Photography [53].
 These results validate our central hypothesis: our data-efficient synthetic training achieves robust physical disentanglement via a low-rank adapter, eluding the limitations of entangled text prompts and rivaling the quality of data-heavy methods.
@@
 Figure 5: Qualitative comparison of physical controls.
-We compare our method against text-based prompting in a T2V backbone (WAN 2.1) and specialized, data-intensive image baselines [Yuan_2024_GenPhoto] and Bokeh Diffusion [fortes2025bokeh].
+We compare our method against text-based prompting in a T2V backbone (WAN 2.1) and specialized, data-intensive image baselines [53] and Bokeh Diffusion [7].
 7 Analysis
@@
 Method: Backbone Intruder Dimension Check.
-We apply the “intruder dimension” analysis from  shuttleworth2024lora.
+We apply the “intruder dimension” analysis from  Shuttleworth et al.
+[34].
 We perform an SVD on the original backbone weight $W_{pre}$ and its final adapted version $W_{lora}=W_{pre}+\Delta W_{lora}$.
@@
 A learning rate of $2{\times}10^{-5}$ with a 100-step warmup is applied.
-Optimization is performed using the AdamW optimizer[loshchilov2017decoupled] with betas $(0.9,0.99)$, weight decay 0.01, and $\varepsilon=10^{-8}$.
+Optimization is performed using the AdamW optimizer[20] with betas $(0.9,0.99)$, weight decay 0.01, and $\varepsilon=10^{-8}$.
 All LoRA and adapter parameters are trained in bfloat16 with activation checkpointing enabled for memory efficiency.